Scores for a particular standardized test are normally distributed with a mean of 80 and a
standard deviation of 14. Find the probability that a randomly chosen score is
i. No greater than 70
ii. At least 95
iii. Between 70 and 95
iv. A student was told that her percentile score on this exam is 72%. Approximately
what is her raw score?

Found 2 solutions by Boreal, Edwin McCravy:
Answer by Boreal(15235)

z=(x-mean)/sd
a. z < (70-80)/14 or -5/7. That probability is 0.2375
b. z> (95-80)/14 or 15/14. That probability is 0.1420
c. The probability for this is 1- the sum of the above two, since this is all that is left. It is 0.6205
Check with 2nd VARS 2 normalcdf (70,95,80,14) ENTER
d. 72nd percentile is where probability=0.7200 VARS 3invnorm (0.72,0,1) yields a z=0.5828
so (x-80)=0.5828*14=8.16
and the score would be 88.16 as a raw score or 88.

Answer by Edwin McCravy(20064)

i. No greater than 70
On your TI-84, 
press 2ND 
press VARS
press 2

Make the screen read: 

    normalcdf
lower:-9999999
upper:70
μ:80
σ:14       
Paste

Scroll down with the down arrow to Paste

Press ENTER

Read on the screen

normalcdf(-9999999,70,80,14)

Press ENTER

Read 0.2375251875 round to 0.2375

--------------------------------------------

ii. At least 95
Same as above except
lower:95
upper:99999999      

Read 0.1419884174 round to 0.1420

--------------------------------------------

iii. Between 70 and 95
Same as above except
lower:70
upper:95

--------------------------------------------

iv. A student was told that her percentile score on this exam is 72%. Approximately
what is her raw score?
On your TI-84, 
press 2ND 
press VARS
press 3

Make the screen read: 

     invnorm
area:0.72
μ:80
σ:14       
Paste

Scroll down with the down arrow to Paste

Press ENTER

Read on the screen

invNorm(0.72,80,14)

Press ENTER

Read 88.15978103  Round to 88.16 or as your teacher tells you.

Edwin